OpenStudy (anonymous):

How do I solve sin^2 x = -2cos x?

OpenStudy (experimentx):

change sin^x in to cos x, you will have quadratic equation in cos x

OpenStudy (inkyvoyd):

Remember, sin^2 x+cos^2 x=1

OpenStudy (inkyvoyd):

solve the above equation that I gave you for sin^2 x, substitute, and do as experimentX has instructed.

OpenStudy (anonymous):

sinx^2+cosx^2=1 \[\sin x =\sqrt{1-\cos^2}\] then u can easily solve it

OpenStudy (anonymous):

So \[\sin ^{2}x=1-\cos ^{2}x\]? And then I replace sin^2x with that?

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

So I wind up with cos^2 x – 2cos x – 1 = 0 and factor it as (cos x – 1)(cos x + 1) then solve?

OpenStudy (inkyvoyd):

No, it doesn't factor that way.

OpenStudy (inkyvoyd):

Add two to both sides to get cos^2 x-2cos x+1=2 then take the square root of both sides (cos x-1)=plus minus (sqrt 2)

OpenStudy (anonymous):

Ah! So confusing.....thank you for you help...I appreciate it

OpenStudy (anonymous):

Also can be left as cosx(cosx-2) =1